OpenStudy (anonymous):

An equation is shown below: 5(2x - 3) = 5 Part A: How many solutions does this equation have? (4 points) Part B: What are the solutions to this equation? Show your work. (6 points)

OpenStudy (anonymous):

sorry, 10x-15=5 10x = 5+15 x= 20/10 x=2

OpenStudy (anonymous):

part A: one solution

syed98 (syedmohammed98):

u can either use the method which ahamed has done or simply divide both the sides with 5 and then solve it :)

OpenStudy (anonymous):

Sasha solved an equation, as shown below: Step 1: 8x = 56 Step 2: x = 56 - 8 Step 3: x = 48 Part A: Is Sasha's solution correct or incorrect? If the solution is incorrect, explain why it is incorrect and show the correct steps to solve the equation. (6 points) Part B: How many solutions does this equation have? (4 points)

OpenStudy (anonymous):

understand, baby?

OpenStudy (anonymous):

A: the solution is incorrect

OpenStudy (anonymous):

A: because we have to divide by 8, not subtract

OpenStudy (anonymous):

B:one solution

OpenStudy (anonymous):

understand, baby?

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

An equation is shown below: 9(3x - 16) + 15 = 6x - 24 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points)

OpenStudy (anonymous):

@Mlink1290 come on, baby. you can do this. USE MY PREVIOUS ANSWER TO GUIDE YOU

OpenStudy (anonymous):

idk!!!!!

syed98 (syedmohammed98):



OpenStudy (anonymous):

so for A I would type A:9(3x - 16) + 15 = 6x - 24 27x - 144 + 5 = 6x - 24 21x = 105 and B: x=answer

syed98 (syedmohammed98):

yupp :)

OpenStudy (anonymous):

thanks!

syed98 (syedmohammed98):

ur welcome :)